# High-Fidelity Data-Driven Dynamics Model for Reinforcement Learning-based Magnetic Control in HL-3 Tokamak

### v1

@@ 4.4.2 Adaptive Weight Adjustment
To unanimously provide multi-task predictions for $25$ variables, in Table 2, the calibration of relative weights assigned to the predictive losses of individual variables could heavily affect the model performance. Mathematically, for a space $\mathcal{V}$ containing $25$ variables, the aggregated loss function given the $44$-dimensional input $\boldsymbol{x}$ shall be a scalar $\mathrm{L}_{\mathcal{V}}\left(\boldsymbol{y}_{\mathcal{V}},\hat{\boldsymbol{y}}_{\mathcal{V}};\boldsymbol{x}\right)$ computed from $25$ individual loss functions $L_{v}\left(y_{v},\hat{y}_{v};\boldsymbol{x}\right)$ weighted by $c_{v}$, where $y_{v}$ and $\hat{y}_{v}$ correspond to actual data and model prediction for the variable $v\in\mathcal{V}$, respectively while $\boldsymbol{y}_{\mathcal{V}}$ and $\hat{\boldsymbol{y}}_{\mathcal{V}}$ denote the concatenation of $y_{v}$ and $\hat{y}_{v}$, $\forall v\in\mathcal{V}$. Due to the discrepancies in distribution and noise among the variables, a straightforward averaging of individual loss functions can adversely impact the model’s accuracy. Meanwhile, manual weight tuning is an arduous and time-intensive effort. Therefore, we utilize an adaptive weight adjustment technique by regarding the loss weight $c_{v}$ as a trainable single variate[45]. To prevent the loss weights from taking negative values during the optimization process, which could lead to negative loss values, we also employ a non-negative regularization constraint [44] to enhance the learning robustness. To sum up, the final loss function for training can be calculated as
$\mathrm{L}_{\mathcal{V}}\left(\boldsymbol{y}_{\mathcal{V}},\hat{\boldsymbol{y}}_{\mathcal{V}};\boldsymbol{x}\right)=\sum_{v\in\mathcal{V}}\left[\frac{1}{2\cdot c_{v}^{2}}\cdot L_{v}\left(y_{v},\hat{y}_{v};\boldsymbol{x}\right)+\ln\left(1+c_{v}^{2}\right)\right].$
@@ 4.4.3 Training Details
Finally, in addition to Fig. 3, Fig. 14 shows the accuracy for predicting all the $18$ coil currents.
Fig. 14: Goodness-of-fit of the dynamics model for coil currents in terms of the coefficient of determination $R^{2}$.
@@ 4.5.1 State, Action, and Reward of RL
In this study, we utilize RL to control the plasma current ($I_{p}$) and six shape and position parameters ($a$, $\kappa$, $\delta_{l}$, $\delta_{u}$, $R$, $Z$), and denote the concatenation of current observations as $\boldsymbol{o}$. To ensure the integrity and stability of HL-3 plasma discharges, we select the actual waveform of shot #6226, an H-mode discharge with ELMs and a plasma current of $300$ kA, as the desired target $\boldsymbol{d}$ for control. As shown in Table 3, we consider a $14$-dimensional state $\boldsymbol{s}$ by amalgamating the observations $\boldsymbol{o}$ and desired target $\boldsymbol{d}$, while the $17$-dimensional action $\boldsymbol{a}$ encompasses voltage commands of $1$ CS coil and $16$ PF coils. To ensure control safety and hardware integrity, we impose operational constraints by stipulating that the dynamic rate of change between two consecutive voltage commands must not exceed $20$ V/ms. In other words, for any coil, any voltage command that changes beyond this limit will be clipped to adhere to the safety constraints.
Table 3: State and action space for RL-based magnetic control.
Name Description
State Space $\mathcal{S}$ Target and currently observed plasma current, and six shape and position parameters
($14$ dimensions)
Action Space $\mathcal{A}$ Voltage commands of CS and $8$ pairs of PF coils
($17$ dimension)
On the other hand, we devise the reward function $r(\boldsymbol{s},\boldsymbol{a})$ from two orthogonal perspectives. First, consistent with [9, 10], the reward function primarily takes account of the element-wise divergence between the current observation $\boldsymbol{o}\triangleq[o^{(1)},\cdots,o^{(7)}]$ and desired target $\boldsymbol{d}\triangleq[d^{(1)},\cdots,d^{(7)}]$. To merge the divergence components into a scalar reward and emphasize those control variables with sub-optimal control results, a smooth-max function [10] is leveraged. Second, owing to the constraints of the power module, the coil current is permitted to cross zero only once. Therefore, contingent on the coil current predicted by the dynamics model, we impose a penalty term to inhibit the agent from exploring control commands that would result in multiple zero crossings of the coil current. To sum up, the reward for a state-action pair $(\boldsymbol{s},\boldsymbol{a})$ is given as
@@
where weights $\boldsymbol{\omega}\triangleq[\omega_{1},\cdots,\omega_{7}]$ in the softmax function provide the (relative) importance of each component and are initialized as $[3,2,2,3,1,2,2]$. The modulation factor $\alpha$, which affects the trade-offs between “easy” and “hard” (to satisfy) components [10], is set as $-1$ in our experiment. Besides, for $17$ coils, the penalty term $p_{i}$ for coil $i$ is designed as $-1$ when the coil current exceeds one zero-crossing while nulls otherwise.
@@ 4.5.2 Dynamics Model-based Training
We utilize the trained dynamics model as the interaction environment, through which the RL agent explores suitable actions to achieve the desired plasma current and shape. Specifically, as the HL-3 tokamak is a preliminarily established device and initiates the very first-round global open campaign this year, a significant number of discharge shots have been scheduled for international joint physical experiments. Consequently, the proposal for RL-based control has to share discharges with other experiments. In addition, given the potential uncertainty in freshly deploying RL-based control on HL-3, the early phase of a discharge, prior to $1500$ ms, is primarily allocated to other physics experiments. Therefore, during training, RL assumes the $1$-kHz control since the $1500$ ms of the discharge, and the initial observations $\boldsymbol{o}$ are synchronized with those of shot #6226, ensuring a smooth transition between PID control and RL control. Furthermore, we take episodic training from $1500$ ms to $2000$ ms of the discharge. Without loss of generality, at each time step $t$, the RL agent determines an action $\boldsymbol{a}_{t}$ according to the state $\boldsymbol{s}_{t}$, while the dynamics model yields next-step plasma current as well as the six shape and position parameters $\boldsymbol{o}_{t+1}$ based on $\boldsymbol{s}_{t}$ and $\boldsymbol{a}_{t}$. Note that given the slow-changing nature of the toroidal magnetic field and loop voltage, these two values are considered invariant and maintain the values same as the handover time of shot #6226. Hence, a reward $r(\boldsymbol{s}_{t},\boldsymbol{a}_{t})$ can be computed as in Equation (1). Finally, the episodic reward is derived from the discounted accumulative reward with a discount factor $\gamma$ set as $0.98$. To strike a balance between the variance and bias in the value function estimation, we employ the Generalized Advantage Estimation (GAE) [78] to compute the advantage function. In addition, other parameters for the PPO are set in a standard way. Specifically, we set the entropy coefficient, the clipping parameter $\epsilon$, and the $\lambda$ parameter for GAE as $0.05$, $0.2$, and $0.95$, respectively. Following this, the PPO algorithm [53] can iteratively update the agent’s actor and critic networks. Fig. 7a illustrates the tracking accuracy of the trained agent toward the target waveform within the simulation setting and suggests that the RL agent, supported by a data-driven dynamics model, has effectively explored and assimilated strategies that are beneficial for achieving discharge control objectives.
@@ 4.5.3 On-Device Deployment
As mentioned earlier, the first global open campaign on HL-3 has attracted a considerable number of international joint experiments, limiting the opportunities for testifying RL-based control. Eventually, in a shared discharge, the handover to RL-based control has to be delayed until other physics experiments involving auxiliary heating are completed. To avoid any potential plasma disturbances caused by shutting down the auxiliary heating at $1500$ ms, which could impact the analysis of experimental results, we have further postponed the start of RL-based control to $1600$ ms. We initially begin with a duration of $15$ ms for RL-based control and then revert to PID control. Afterward, we gradually extend the duration upon each successful execution, and ultimately, after $6$ effective shots, the experiment ends with a control duration of $100$ ms, with the last remaining shot attempting to implement a zero-shot control towards changed triangularity targets. Notably, to ensure stability incurred during these mode switches, we establish a $10$-ms transition period, during which a weighted sum of commands from both modes is undertaken and the weight of superseding control mode linearly increases. The limited opportunities hinder the on-device experiment for a longer RL-based takeover in practice.
@@ 4.5.4 Post-Experiment Analyses
As illustrated in Fig. 7b and Fig. 9a, the post-experiment replay for shot #$6691$ and shot #$6698$ is calculated using the dynamics model. Specifically, the dynamics model starts to evolve according to the input variables in Table 2, set as practical values at the handover time, as well as the practically executed coil voltage commands, which are generated by RL during on-device deployment. In other words, such a post-experiment replay represents the autoregressive output of the dynamics model, and its closeness to the EFITNN-based diagnostic measurement contributes to verifying the accuracy of the dynamic model.
On the other hand, as shown in Fig. 4d and Fig. 9b, to visually display the prediction and control effect on the plasma shape, we calculate the four extreme points of the cross-section based on the six shape parameters obtained from the EFITNN computation. The plasma boundary is then approximated as the connection of four-quarter ellipses [80].
@@ Acknowledgments
The authors would like to thank the entire HL-3 team for providing experimental data and engineering support for the verification experiments in the HL-3 tokamak. This work was in part supported by the National Nature Science Foundation of China under Grant No. U21A20440.
@@
Data availability
Data sets generated during the current study are available from the corresponding author on request.
Code availability
Codes used during the current study are available from the corresponding author on request.

### v2

@@ 4.4.2 Adaptive Weight Adjustment
To unanimously provide multi-task predictions for $25$ variables, in Supplementary Table 2, the calibration of relative weights assigned to the predictive losses of individual variables could heavily affect the model performance. Mathematically, for a space $\mathcal{V}$ containing $25$ variables, the aggregated loss function given the $44$-dimensional input $\boldsymbol{x}$ shall be a scalar $\mathrm{L}_{\mathcal{V}}\left(\boldsymbol{y}_{\mathcal{V}},\hat{\boldsymbol{y}}_{\mathcal{V}};\boldsymbol{x}\right)$ computed from $25$ individual loss functions $L_{v}\left(y_{v},\hat{y}_{v};\boldsymbol{x}\right)$ weighted by $c_{v}$, where $y_{v}$ and $\hat{y}_{v}$ correspond to actual data and model prediction for the variable $v\in\mathcal{V}$, respectively while $\boldsymbol{y}_{\mathcal{V}}$ and $\hat{\boldsymbol{y}}_{\mathcal{V}}$ denote the concatenation of $y_{v}$ and $\hat{y}_{v}$, $\forall v\in\mathcal{V}$. Due to the discrepancies in distribution and noise among the variables, a straightforward averaging of individual loss functions can adversely impact the model’s accuracy. Meanwhile, manual weight tuning is an arduous and time-intensive effort. Therefore, we utilize an adaptive weight adjustment technique by regarding the loss weight $c_{v}$ as a trainable single variate[45]. To prevent the loss weights from taking negative values during the optimization process, which could lead to negative loss values, we also employ a non-negative regularization constraint [44] to enhance the learning robustness. To sum up, the final loss function for training can be calculated as
$\mathrm{L}_{\mathcal{V}}\left(\boldsymbol{y}_{\mathcal{V}},\hat{\boldsymbol{y}}_{\mathcal{V}};\boldsymbol{x}\right)=\sum_{v\in\mathcal{V}}\left[\frac{1}{2\cdot c_{v}^{2}}\cdot L_{v}\left(y_{v},\hat{y}_{v};\boldsymbol{x}\right)+\ln\left(1+c_{v}^{2}\right)\right].$ (1)
@@ 4.4.3 Training Details
Finally, in addition to Fig. 3, Supplementary Figure 5 shows the accuracy for predicting all the $18$ coil currents.
@@ 4.5.1 State, Action, and Reward of RL
In this study, we utilize RL to control the plasma current ($I_{\mathrm{p}}$) and six shape and position parameters ($a$, $\kappa$, $\delta_{\mathrm{l}}$, $\delta_{\mathrm{u}}$, $R$, $Z$), and denote the concatenation of current observations as $\mathbf{o}$. To ensure the integrity and stability of HL-3 plasma discharges, we select the actual waveform of shot #6226, an H-mode discharge with ELMs and a plasma current of $300$ kA, as the desired target $\mathbf{d}$ for control. As shown in Supplementary Table 3, we consider a $14$-dimensional state $\mathbf{s}$ by amalgamating the observations $\mathbf{o}$ and desired target $\mathbf{d}$, while the $17$-dimensional action $\mathbf{a}$ encompasses voltage commands of $1$ CS coil and $16$ PF coils. To ensure control safety and hardware integrity, we impose operational constraints by stipulating that the dynamic rate of change between two consecutive voltage commands must not exceed $20~\mathrm{V}~\mathrm{ms}^{-1}$. In other words, for any coil, any voltage command that changes beyond this limit will be clipped to adhere to the safety constraints.
Typically, RL relies on explicit targets for reward function design. In fusion experiments, however, diverse targets from varied discharge templates complicate reward engineering and can undermine model efficacy. Therefore, we devise the reward function $r(\boldsymbol{s},\boldsymbol{a})$ from two orthogonal perspectives. First, consistent with [9, 10], the reward function primarily takes account of the element-wise divergence between the current observation $\boldsymbol{o}\triangleq[o^{(1)},\cdots,o^{(7)}]$ and desired target $\boldsymbol{d}\triangleq[d^{(1)},\cdots,d^{(7)}]$. To merge the divergence components into a scalar reward and emphasize those control variables with sub-optimal control results, a smooth-max function [10] is leveraged. Second, owing to the constraints of the power module, the coil current is permitted to cross zero only once. Therefore, contingent on the coil current predicted by the dynamics model, we impose a penalty term to inhibit the agent from exploring control commands that would result in multiple zero crossings of the coil current. To sum up, the reward for a state-action pair $(\boldsymbol{s},\boldsymbol{a})$ is given as
@@
where weights $\boldsymbol{\omega}\triangleq[\omega_{1},\cdots,\omega_{7}]$ in the softmax function provide the (relative) importance of each component and are initialized as $[3,2,2,3,1,2,2]$. The modulation factor $\alpha$, which affects the trade-offs between “easy” and “hard” (to satisfy) components [10], is set as 1-1 in our experiment. Besides, for $17$ coils, the penalty term $p_{i}$ for coil $i$ is designed as 1-1 when the coil current exceeds one zero-crossing while nulls otherwise.
@@ 4.5.2 Dynamics Model-based Training
We utilize the trained dynamics model as the interaction environment, through which the RL agent explores suitable actions to achieve the desired plasma current and shape. Specifically, as the HL-3 tokamak is a preliminarily established device and initiates the global open campaigns in 2024 and 2025, a significant number of discharge shots have been scheduled for international joint physical experiments. Consequently, the proposal for RL-based control has to share discharges with other experiments. The choice of shot #6226 for target tracking is also motivated by the potential to maximize RL’s takeover opportunities, considering the frequent use of similar shot templates in planned experiments on the HL-3 tokamak. As highlighted in Fig. 9, the promising outcomes underscore the inherent generalization capability, which suggests robust applicability to a variety of other discharge templates. In addition, given the potential uncertainty in freshly deploying RL-based control on HL-3, the early phase of a discharge, prior to $1500$ ms, is primarily allocated to other physics experiments. Therefore, during training, RL assumes the $1$-kHz control since the $1500$ ms of the discharge, and the initial observations $\boldsymbol{o}$ are synchronized with those of shot #6226, ensuring a smooth transition between PID control and RL control. Furthermore, we take episodic training from $1500$ ms to $2000$ ms of the discharge. Without loss of generality, at each time step $t$, the RL agent determines an action $\mathbf{a}_{t}$ according to the state $\mathbf{s}_{t}$, while the dynamics model yields next-step plasma current as well as the six shape and position parameters $\mathbf{o}_{t+1}$ based on $\mathbf{s}_{t}$ and $\mathbf{a}_{t}$. Note that given the slow-changing nature of the toroidal magnetic field and loop voltage, these two values are considered invariant and maintain the values same as the handover time of shot #6226. Hence, a reward $r(\mathbf{s}_{t},\mathbf{a}_{t})$ can be computed as in Equation (2). Finally, the episodic reward is derived from the discounted accumulative reward with a discount factor $\gamma$ set as $0.98$. To strike a balance between the variance and bias in the value function estimation, we employ the Generalized Advantage Estimation (GAE) [69] to compute the advantage function. In addition, other parameters for the PPO are set in a standard way. Specifically, we set the entropy coefficient, the clipping parameter $\epsilon$, and the $\lambda$ parameter for GAE as $0.05$, $0.2$, and $0.95$, respectively. Following this, the PPO algorithm [53] can iteratively update the agent’s actor and critic networks. Fig. 7a illustrates the tracking accuracy of the trained agent toward the target waveform within the simulation setting and suggests that the RL agent, supported by a data-driven dynamics model, has effectively explored and assimilated strategies that are beneficial for achieving discharge control objectives.
The aforementioned single-template training necessitates retraining the model with a new target for changed operational scenarios (e.g., targeting a $700$ kA plasma current). Despite this requirement for retraining, the process is efficient. Systematic evaluations across different GPU platforms (Supplementary Figure 9 and Supplementary Table 4) substantiate this efficiency. The PPO-based controller converges within $500$ training episodes ($25,000$ interaction steps). Utilizing a standard configuration (PyTorch 2.1.0, CUDA 12.2, batch size $512$), training on an NVIDIA GeForce RTX 4090 completes in just $22$ minutes. Such rapid training underscores the framework’s practical adaptability, even with the need for template-specific retraining.
@@ 4.5.3 On-Device Deployment
As mentioned earlier, the open campaigns on HL-3 have attracted a considerable number of international joint experiments, limiting the opportunities for testifying RL-based control. Eventually, in a shared discharge, the handover to RL-based control has to be delayed until other physics experiments involving auxiliary heating are completed. To avoid any potential plasma disturbances caused by shutting down the auxiliary heating at $1500$ ms, which could impact the analysis of experimental results, we have further postponed the start of RL-based control to $1600$ ms. In other words, we activate the RL control at $1600$ ms and revert to the PID control after a predefined window. To ensure stability incurred during these mode switches, we establish a transition period, during which a weighted sum of commands from both modes is undertaken and the weight of superseding control mode linearly increases.
Our initial RL control experiments are conducted during the #6691-#6698 shot series on HL-3. Due to the coincidence with the development of the fast vertical position control system, the experiments undergo severely unstable discharge conditions, with only three effective discharges (i.e., shots #6691 (Supplementary Figure 6a) , #6693 (Supplementary Figure 7) and #6698 (Supplementary Figure 6b)) sustaining the flat-top phase beyond $1600$ ms. Furthermore, to ensure operational safety, the RL control window is limited to $100$ ms, with $10$ ms transition periods at its start and end. Unfortunately, the conventional PID and fast vertical position control still exhibits limited capability to suppress emerging voltage variations after reverting from RL to PID control, frequently resulting in plasma disruptions. More recently, following the maturation of the fast vertical position control system for routine operations, we replicate these experiments during the #12777-#12781 shot series, by adopting the same actor model. This second campaign yields three effective RL-controlled discharges: shots #12777 ($100$ ms), #12779 ($260$ ms), and #12781 ($400$ ms). For these shots, $30$ ms transition periods are implemented at the start and end. All three effective discharges are completed with no disruptive events. Fig. 7b illustrates the complete discharge for shot #12781. The complete discharges for the remaining five successful RL-controlled shots (from both experimental campaigns) are presented in Supplementary Figures 6, 7, and 8. To quantitatively assess the performance differences between RL–based control and fixed voltage control, we conduct a comparative experiment using shot #12777. For shot #12777, it adopts an RL control solely until $1700$ ms without the last $30$-ms transition period, while from $1700$ ms to $1730$ ms, the control voltages are deliberately fixed at their $1700$ ms values. As depicted in Supplementary Figure 8, the resulting oscillations in plasma current and shape confirm that, without closed-loop control, plasma stability is significantly compromised, demonstrating the critical need for an effective closed-loop control. Supplementary Figure 10 illustrates the plasma current for the seven shots from these campaigns in which RL control is not executed.
@@ 4.5.4 Post-Experiment Analyses
As illustrated in Fig. 7b and Fig. 9a, the post-experiment replay for shot #12781 and shot #6698 is calculated using the dynamics model. Specifically, the dynamics model starts to evolve according to the input variables in Supplementary Table 2, set as practical values at the handover time, as well as the practically executed coil voltage commands, which are generated by RL during on-device deployment. In other words, such a post-experiment replay represents the autoregressive output of the dynamics model, and its closeness to the EFITNN-based diagnostic measurement contributes to verifying the accuracy of the dynamic model.
On the other hand, as shown in Fig. 4d, to visually display the prediction and control effect on the plasma shape and position, we calculate the four extreme points of the cross-section based on the six shape and position parameters obtained from the EFITNN computation. The plasma boundary is then approximated as the connection of four-quarter ellipses [71].
To better understand the control logic of the RL controller, we perform a more in-depth analysis of the experimental results from shot #$12781$. As illustrated in Fig. 7b, while the controller demonstrates high-fidelity tracking of the primary shape parameters ($a$, $\kappa$, $\delta_{\mathrm{u}}$ and $\delta_{\mathrm{l}}$), the positional parameters ($R$ and $Z$) exhibit persistent steady-state deviations from their targets. This differential control efficacy is primarily attributed to three factors: (1) the inherent strong coupling among the shape-defining parameters ($a$, $\kappa$, $\delta_{\mathrm{l}}$ and $\delta_{\mathrm{u}}$), which are co-regulated by the PF coils; (2) the non-negligible influence of non-electromagnetic experimental disturbances, such as gas puffing and auxiliary heating, which could significantly affect the evolution of $R$ and $Z$; and (3) an implicit control priority mechanism within the RL policy – the policy prioritizes the correction of $\kappa$, $\delta_{\mathrm{l}}$, $\delta_{\mathrm{u}}$, and $R$ due to the larger deviations therein. In summary, compared to the standard PID controller, the RL controller obtains similar precision in regulating $R$ and $Z$, but significantly extends this capability by simultaneously providing closed-loop control over minor radius $a$, upper and lower triangularity $\delta_{\mathrm{u}}$, $\delta_{\mathrm{l}}$, and elongation $\kappa$.
Data availability
Data sets generated during the current study are available from the corresponding author on reasonable request.
Code availability
Codes used during the current study are available from the corresponding author on reasonable request.
@@ Acknowledgments
The authors would like to thank the entire HL-3 team for providing experimental data and engineering support for the verification experiments in the HL-3 tokamak. This work was supported in part by the National Key Research and Development Program of China under Grant 2024YFE03020001, the Zhejiang Provincial Natural Science Foundation of China under Grant LR23F010005, the National Nature Science Foundation of China under Grant No. U21A20440 and No. 12405253.